# Bayesian Deep Learning for Discrete Choice

### v1

@@ 4.1 Approximation of the Posterior for Bayesian Deep Learning
There are multiple methods to approximate the posterior distribution $p(\bm{\Theta}|\bm{y},\bm{X},\bm{Q})$ in Bayesian deep learning [13]. Some of these methods include, but are not limited to posterior sampling [e.g. [15]] and Laplace and variational approximations [e.g. [26]]. Ultimately, the primary goal is to approximate or sample from the posterior distribution in regions of high density while capturing a set of parameters that induce an informative distribution–that captures epistemic and aleatoric uncertainty–over function predictions, $f(\bm{X},\bm{Q};\bm{\Theta})$. In discrete choice, this function corresponds to the latent representative utility $\bm{v}$, as defined in previous sections.
In other words, if a finite number of parameter sets $\bm{\Theta}$ are used to approximate the posterior, each $\bm{\Theta}$ should have a high posterior probability while collectively capturing epistemic and aleatoric uncertainty. This is achieved by having estimates of $f(\bm{X},\bm{Q};\bm{\Theta})$ that vary in regions of the input space with scarce observations and converge in regions where data is abundant [20] and aleatoric uncertainty is low. A good approximation to the posterior distribution will ensure a good representation of uncertainty, which is illustrated in figure 4
@@ 4.1.1 Deep Ensembles
Deep ensembles are models that have empirically demonstrated improvements in accuracy over single deep learning models [25]. A deep ensemble is essentially a collection of deep learning models that are used collectively to make predictions. In the past, deep ensembles have been considered alternatives to Bayesian deep learning; however, Wilson and Izmailov argued that they are simply another way to approximate the posterior [20].
Deep ensembles can be trained in parallel using a single dataset and model architecture. Lakshminarayanan et al. [25] developed an adversarial algorithm to train each model independently. However, the adversarial component can be omitted, and a deep ensemble can be constructed by optimizing the posterior density multiple times from different random initializations in the weight space. They propose that the predictions of each model in the ensemble are weighted equally—equivalently assuming that each weight configuration has the same a posteriori probability. Consequently, the predictions from a deep ensemble are given by:
@@ 4.1.2 Stochastic Weight Average (SWA) and SWA-Gaussian (SWAG)
Stochastic Weight Averaging (SWA) is a computationally efficient way to deal with unstable parameter estimates. SWA offers better generalization and more stable solutions than those obtained from standard training methods while potentially providing better behavioral insights. The SWA approach averages the parameter iterates from stochastic gradient descent during the learning process. The algorithm introduces minimal computational and memory overhead since the parameter average is computed as a running average every $c$ gradient updates. A simplified version of the learning procedure, presented in [23], is shown below:
Algorithm 1 Stochastic Weight Averaging [23]
@@
2:$\bm{\Theta}\leftarrow\tilde{\bm{\Theta}}$triangleright Initialize parameters with $\tilde{\bm{\Theta}}$
3:$\bm{\Theta}_{SWA}\leftarrow\tilde{\bm{\Theta}}$triangleright Initialize SWA parameters
4:$\eta_{models}\leftarrow 0$triangleright Initialize number of models in average
@@
6:     $\bm{\Theta}\leftarrow\bm{\Theta}-\alpha\nabla L(\bm{\Theta})$triangleright Stochastic gradient update
@@
8:         $\bm{\Theta}_{SWA}\leftarrow\frac{\bm{\Theta}_{SWA}\cdot\eta_{models}+\bm{\Theta}}{\eta_{models}+1}$triangleright Compute model average
9:         $\eta_{models}\leftarrow\eta_{models}+1$triangleright Increment the number of models in the average
SWAG has a very similar implementation to SWA but differs in that it also computes the running parameters’ variance and uses these statistics to approximate the posterior with a high-dimensional multivariate normal distribution. In that sense, SWA only provides a set of parameters in a flatter region of the parameter space that has the potential for better generalization [23] but remains a single point estimate, while SWAG gives a true approximation to the parameter posterior that can be used for Bayesian inference.
@@ 4.1.3 Stochastic Gradient Langevin Dynamics
With this gradient update structure, $\bm{\Theta}_{t}$ approaches samples from the posterior $p(\bm{\Theta}\mid q_{ti},x_{ti},y_{ti})$, as these updates approximate Langevin dynamics, which converge to the posterior distribution [15]. In contrast with SWAG, SGLD does not approximate the posterior mode using a Gaussian distribution and does not require additional post-estimation sampling.
@@ 5 A General Deep Learning Discrete Choice Model Architecture
Our model architecture builds on some of the foundations we established in [10]. It consists of four distinct components:
@@
    An embedding layer for shared input variables across alternatives (as presented in[18] for better interpretability).
@@
where $\bm{q}_{ij}$ is the embedding representation for alternative $j$ and individual characteristics $\bm{q}_{i}$. The BatchNorm layers are not affine, meaning that the output is not rescaled and relocated. Instead, parameters $\sigma_{IIA}$ and $\sigma_{nonIIA}$ are introduced to scale the outputs from those components, but we purposely exclude the relocation parameters as the location of the latent utilities $v_{ij}$ with respect to the base alternative is already determined by the alternative specific constant $\alpha_{j}$.
The linear component of the model, $\alpha_{j}+\bm{x}_{ij}\bm{\beta}+\bm{q}_{ij}\bm{\gamma}$, is used to represent the part of the model informed by expert knowledge and behavioral assumptions. However, in practice, this component can have any other form and include parameter interactions and non-linear parameterized components. Creating embeddings for $\bm{q}_{ij}$ is necessary to ensure identifiability of the linear model, similar to the approach proposed in [18]. The main difference is that we allow individual embeddings to interact with alternative-specific attributes in the non-linear IIA block. However, alternatively, the informed component could be formulated with $\bm{\gamma}_{J}=\bm{0}$, or adapted to incorporate other assumptions relevant to each case study.
The non-linear component consists of an IIA block, $\text{BatchNorm}[f_{\Theta_{f}}(\bm{x}_{ij},\bm{q}_{ij})]$ parametrized by $\Theta_{f}$, and a non-IIA block, $\text{BatchNorm}[g_{\Theta_{g}}(\bm{x}_{i},\bm{q}_{i})_{j}]$ parametrized by $\Theta_{g}$. Batch normalization stabilizes the model parameters by computing batch statistics to normalize outputs across the batch, preventing internal covariate shift [27]. Additionally, batch normalization has been shown to accelerate training and improve performance [27]. In this model, the BatchNorm layers do not have any learnable parameters, and the outputs are rescaled (by $\sigma_{IIA}$ or $\sigma_{nonIIA}$) but not shifted. Its inclusion in our neural network can be interpreted as a mechanism to determine the general location of the representative utilities while providing a straightforward–interpretable–way to impose priors on the scale of the non-linear components (by penalizing the scales $\sigma_{IIA}$ or $\sigma_{nonIIA}$ instead of relying on standard $\ell_{2}$ or $\ell_{1}$ regularization, which penalizes the norms of all neural network weights).
@@
To further ensure that our model defaults to the behaviorally informed hypothesis when the data does not support more complex explanations, we propose a two-step training procedure. First, we optimize the cost function (posterior) while keeping the nonlinear layers frozen until convergence. Then, we unfreeze the nonlinear layers and begin the SGLD sampling process. This two-step learning procedure is designed to guide the sampler toward regions of the parameter space where the behaviorally informed coefficients have high magnitudes. It also ensures that the model only learns more complex representations when they provide improvements over the knowledge-informed baseline.
@@ 6 Monte Carlo Study for Empirical Coverage
To estimate the performance of our model, we employ a Monte Carlo simulation study, similar to the one presented in [28] for Latent Class discrete choice models. Here, we assume that the latent utility function follows the equation:
@@ 7 Revealed and Stated Preference Case Studies
In addition to the simulation study, we consider two case studies: one with revealed preference data and another with stated preference data. We use the revealed preference dataset constructed in [29] for multinomial mode choice in NYC. For the stated preference data, we use the popular and publicly available [^4] Available in Apollo in R as the apollo_swissRouteChoiceData. dataset from a binary train choice experiment in Switzerland [30].
@@ 7.1 Mode Choice in New York City
High income Indicator variable for high income level ($>100$k USD per year). $0.32$
@@ 8.1 Simulation Study
The empirical coverages for the MRS, illustrated in Figures 6 to 8 and summarized in Table 3, show that: i) our model achieves empirical coverage consistent with a 95% credibility level, regardless of dataset size; ii) the fully connected neural network fails to achieve adequate coverage for the smaller datasets but attains coverage close to 95% for the larger datasets; and iii) the model that ignores nonlinearities (the conditional logit model) achieves coverage close to 95% only for MRS associated with alternative attributes that have linear or near-linear relationships with the latent utilities, as we will show next [^6] Alternative attributes $x_{1}$ and $x_{2}$ have a linear relationship with the latent utilities, so the first MRS should be accurately estimated by a linear model..
@@ 8.2 Mode Choice in New York City
For the revealed preference case study using New York City multinomial mode choice data, we estimated values of travel time savings (VOTT) for train and car, as well as the out-of-sample accuracy of our model and a fully connected neural network. Standard discrete choice models, graph neural networks, and hierarchical Gaussian process models were previously estimated in [29] and [10].
@@
SOTA - 78.45  [^7] From the Skip-GNN presented by us in [10].

### v2

@@ 4.1 Approximation of the Posterior for Bayesian Deep Learning
There are multiple methods to approximate the posterior distribution $p(\bm{\Theta}|\bm{y},\bm{X},\bm{Q})$ in Bayesian deep learning [papamarkou2024position]. Some of these methods include, but are not limited to posterior sampling [e.g. [welling2011bayesian]] and Laplace and variational approximations [e.g. [10.1162/neco.1992.4.3.415]]. Ultimately, the primary goal is to approximate or sample from the posterior distribution in regions of high density while capturing a set of parameters that induce an informative distribution–that captures epistemic and aleatoric uncertainty–over function predictions, $f(\bm{X},\bm{Q};\bm{\Theta})$. In discrete choice, this function corresponds to the latent representative utility $\bm{v}$, as defined in previous sections.
In other words, if a finite number of parameter sets $\bm{\Theta}$ are used to approximate the posterior, each $\bm{\Theta}$ should have a high posterior probability while collectively capturing epistemic and aleatoric uncertainty. This is achieved by having estimates of $f(\bm{X},\bm{Q};\bm{\Theta})$ that vary in regions of the input space with scarce observations and converge in regions where data is abundant [Wilson2020] and aleatoric uncertainty is low. A good approximation to the posterior distribution will ensure a good representation of uncertainty, which is illustrated in figure 4
@@ 4.1.1 Deep Ensembles
Deep ensembles are models that have empirically demonstrated improvements in accuracy over single deep learning models [Lakshminarayanan2017]. A deep ensemble is essentially a collection of deep learning models that are used collectively to make predictions. In the past, deep ensembles have been considered alternatives to Bayesian deep learning; however, Wilson and Izmailov argued that they are simply another way to approximate the posterior [Wilson2020].
Deep ensembles can be trained in parallel using a single dataset and model architecture. Lakshminarayanan et al. [Lakshminarayanan2017] developed an adversarial algorithm to train each model independently. However, the adversarial component can be omitted, and a deep ensemble can be constructed by optimizing the posterior density multiple times from different random initializations in the weight space. They propose that the predictions of each model in the ensemble are weighted equally—equivalently assuming that each weight configuration has the same a posteriori probability. Consequently, the predictions from a deep ensemble are given by:
@@ 4.1.2 Stochastic Weight Average (SWA) and SWA-Gaussian (SWAG)
Stochastic Weight Averaging (SWA) is a computationally efficient way to deal with unstable parameter estimates. SWA offers better generalization and more stable solutions than those obtained from standard training methods while potentially providing better behavioral insights. The SWA approach averages the parameter iterates from stochastic gradient descent during the learning process. The algorithm introduces minimal computational and memory overhead since the parameter average is computed as a running average every $c$ gradient updates. A simplified version of the learning procedure, presented in [izmailov2018averaging], is shown below:
Algorithm 1 Stochastic Weight Averaging [izmailov2018averaging]
@@
[…]
@@
6:  $\bm{\Theta}\leftarrow\bm{\Theta}-\alpha\nabla L(\bm{\Theta})$ $\triangleright$ Stochastic gradient update
@@
8:   $\bm{\Theta}_{SWA}\leftarrow\frac{\bm{\Theta}_{SWA}\cdot\eta_{models}+\bm{\Theta}}{\eta_{models}+1}$ $\triangleright$ Compute model average
9:   $\eta_{models}\leftarrow\eta_{models}+1$ $\triangleright$ Increment the number of models in the average
SWAG has a very similar implementation to SWA but differs in that it also computes the running parameters’ variance and uses these statistics to approximate the posterior with a high-dimensional multivariate normal distribution. In that sense, SWA only provides a set of parameters in a flatter region of the parameter space that has the potential for better generalization [izmailov2018averaging] but remains a single point estimate, while SWAG gives a true approximation to the parameter posterior that can be used for Bayesian inference.
@@ 4.1.3 Stochastic Gradient Langevin Dynamics
With this gradient update structure, $\bm{\Theta}_{t}$ approaches samples from the posterior $p(\bm{\Theta}\mid q_{ti},x_{ti},y_{ti})$, as these updates approximate Langevin dynamics, which converge to the posterior distribution [welling2011bayesian]. In contrast with SWAG, SGLD does not approximate the posterior mode using a Gaussian distribution and does not require additional post-estimation sampling.
@@ 5 A General Deep Learning Discrete Choice Model Architecture
Our model architecture builds on some of the foundations we established in [2025arXiv250309786V]. It consists of four distinct components:
@@
    An embedding layer for shared input variables across alternatives (as presented in[arkoudi2023combining] for better interpretability).
@@
where $\bm{q}_{ij}$ is the embedding representation for alternative $j$ and individual characteristics $\bm{q}_{i}$. The BatchNorm layers are not affine, meaning that the output is not rescaled and relocated. Instead, parameters $\sigma_{IIA}$ and $\sigma_{nonIIA}$ are introduced to scale the outputs from those components, but we purposely exclude the relocation parameters as the location of the latent utilities $v_{ij}$ with respect to the base alternative is already determined by the alternative specific constant $\alpha_{j}$. The BatchNorm layers also ensure that the nonlinear component of our model does not overtake the knowledge-informed part of the model, which may be as simple as being linear in attributes and parameters.
This architecture resembles that proposed by Wang et al. [wang2021theory] in that it combines a knowledge-informed component with a deep learning model capable of automatic feature learning. However, their architecture (i) does not fix the location of the utilities, (ii) does not include separate IIA and non-IIA nonlinear components, and (iii) introduces a weighting parameter $\lambda$ for the linear combination of components, but does not relate it directly to their relative scale, since the components are not explicitly scaled (which could hinder convergence). Our model incorporates these three points.
The linear component of the model, $\alpha_{j}+\bm{x}_{ij}\bm{\beta}+\bm{q}_{ij}\bm{\gamma}$, is used to represent the part of the model informed by expert knowledge and behavioral assumptions. However, in practice, this component can have any other form and include parameter interactions and non-linear parameterized components. Creating embeddings for $\bm{q}_{ij}$ is necessary to ensure identifiability of the linear model, similar to the approach proposed in [arkoudi2023combining]. The main difference is that we allow individual embeddings to interact with alternative-specific attributes in the non-linear IIA block. However, alternatively, the informed component could be formulated with $\bm{\gamma}_{J}=\bm{0}$, or adapted to incorporate other assumptions relevant to each case study.
The non-linear component consists of an IIA block, $\text{BatchNorm}[f_{\Theta_{f}}(\bm{x}_{ij},\bm{q}_{ij})]$ parametrized by $\Theta_{f}$, and a non-IIA block, $\text{BatchNorm}[g_{\Theta_{g}}(\bm{x}_{i},\bm{q}_{i})_{j}]$ parametrized by $\Theta_{g}$. Batch normalization stabilizes the model parameters by computing batch statistics to normalize outputs across the batch, preventing internal covariate shift [ioffe2015batch]. Additionally, batch normalization has been shown to accelerate training and improve performance [ioffe2015batch]. In this model, the BatchNorm layers do not have any learnable parameters, and the outputs are rescaled (by $\sigma_{IIA}$ or $\sigma_{nonIIA}$) but not shifted. Its inclusion in our neural network can be interpreted as a mechanism to determine the general location of the representative utilities while providing a straightforward–interpretable–way to impose priors on the scale of the non-linear components (by penalizing the scales $\sigma_{IIA}$ or $\sigma_{nonIIA}$ instead of relying on standard $\ell_{2}$ or $\ell_{1}$ regularization, which penalizes the norms of all neural network weights).
@@
5.1 Interpretability and Identifiability
It is well known that deep learning models are not identifiable. However, the structure of our model, coupled with a two-step learning procedure, stabilizes parameter estimates around a convergence mode in a region of the parameter space where simple hypotheses attain large log-posterior values (relative to other regions). Our proposed two-step learning process is depicted in Figures 7 and 7.
Figure 6: Step 1: Convergence toward a region (green) of the parameter space where simple hypotheses attain large log-posterior values.
Figure 7: Step 2: Sampling (black dots) close to a convergence mode within a region where simple hypotheses attain large log-posterior values.
@@ 5.1 Interpretability and Identifiability
Our learning procedure ensures that the model defaults to the behaviorally informed hypothesis when the data do not support more complex explanations. First, we optimize the objective (log-posterior) while keeping the nonlinear layers frozen until convergence (Figure 7). We then unfreeze the nonlinear layers and initiate SGLD sampling (Figure 7). This procedure is designed to guide the sampler toward regions of the parameter space in which the behaviorally informed coefficients retain large magnitudes and simple hypotheses attain high log-posterior values. It also ensures that the model learns more complex representations only when they improve upon the knowledge-informed baseline. Finally, by biasing sampling toward posterior modes associated with simple hypotheses, the procedure can improve MCMC mixing by avoiding exploration of modes that are poorly supported or behaviorally implausible.
5.2 Uncertainty Propagation from Model Parameters to Behavioral Variables
In order for approximate Bayesian inference to be valid for behavioral variables, it must be shown how to propagate uncertainty in the model parameters propagates to uncertainty in functions of those parameters. Important behavioral variables are functions of (i) representative utilities, (ii) marginal utilities, and (iii) choice probabilities [wang2020deep]. Since marginal rates of substitution are functions of marginal utilities, we focus on the posterior of the marginal-utility effect for attribute $k$ and alternative $j$ at observation $i$.
Let $\theta\equiv(\beta,\Theta_{f},\Theta_{g})$ denote all model parameters and define the marginal-utility effect as the deterministic function:
$m_{ijk}(\theta)\;\equiv\;\frac{\partial u_{ij}}{\partial(x_{ij})_{k}}\!\left(\bm{x}_{ij},\bm{q}_{ij};\theta\right)=\beta_{k}+\sigma_{\text{IIA}}\,\frac{\partial f_{\Theta_{f}}(\bm{x}_{ij},\bm{q}_{ij})}{\partial[\bm{x}_{ij}]_{k}}+\sigma_{\text{non-IIA}}\,\frac{\partial g_{\Theta_{g}}\!\left((\bm{x}_{i},\bm{q}_{i})_{j}\right)}{\partial[\bm{x}_{ij}]_{k}}.$ (14)
Here, $\sigma_{\text{IIA}}$ and $\sigma_{\text{non-IIA}}$ are fixed because they are treated as hyperparameters during learning. A point estimate (e.g., a local MAP within a convergence mode) is obtained by evaluating the above expression at $\hat{\theta}$.
A Bayesian analysis, however, requires the posterior distribution of $m_{ijk}$ induced by the posterior over parameters. Since $m_{ijk}$ is a deterministic function of $\theta$, its posterior is the pushforward of $p(\theta\mid D)$ through $m_{ijk}(\cdot)$:
$\mathbb{P}\!\left(m_{ijk}\in B\mid\bm{x}_{ij},\bm{q}_{ij},D\right)=\int\mathrm{1}\left(m_{ijk}(\theta)\in B\right)p(\theta\mid D)\,d\theta,$ (15)
With SGLD (and MCMC approximations in general), we obtain approximate posterior samples $\{\theta_{t}\}_{t=1}^{T}$ and propagate uncertainty by evaluating the effect at each draw,
$m^{(t)}_{ijk}\;=\;m_{ijk}(\theta_{t}),\qquad\theta_{t}\sim p(\theta\mid D),$ (16)
which yields the empirical approximation
$\mathbb{P}\!\left(m_{ijk}\in B\mid\bm{x}_{ij},\bm{q}_{ij},D\right)\;\approx\;\frac{1}{T}\sum_{t=1}^{T}\mathrm{1}\left(m_{ijk}(\theta^{(t)})\in B\right).$ (17)
Posterior moments can be computed by Monte Carlo integration, e.g.,
$\mathbb{E}\!\left[m_{ijk}\mid\bm{x}_{ij},\bm{q}_{ij},D\right]\;\approx\;\frac{1}{T}\sum_{t=1}^{T}m^{(t)}_{ijk},\qquad\mathrm{Var}\!\left(m_{ijk}\mid\bm{x}_{ij},\bm{q}_{ij},D\right)\;\approx\;\frac{1}{T-1}\sum_{t=1}^{T}\left(m^{(t)}_{ijk}-\bar{m}_{ijk}\right)^{2},$ (18)
where $\bar{m}_{ijk}=\mathbb{E}\!\left[m_{ijk}\mid\bm{x}_{ij},\bm{q}_{ij},D\right]\;$. Finally, the $(1-\alpha)$ credible interval for the marginal-utility effect is computed using posterior quantiles:
$\mathrm{CI}_{1-\alpha}\!\left(m_{ijk}\right)\;\approx\;\Big[\,Q_{\alpha/2}\!\left(\{m^{(t)}_{ijk}\}_{t=1}^{T}\right),\;Q_{1-\alpha/2}\!\left(\{m^{(t)}_{ijk}\}_{t=1}^{T}\right)\,\Big].$ (19)
The same procedure applies to any behavioral quantity that is a deterministic function of parameters and covariates, including marginal rates of substitution (as functions of marginal utilities) and quantities derived from choice probabilities.
5.3 Epistemic and Aleatoric Uncertainty
In this study, we focus on uncertainty representation and leave the disentanglement of epistemic and aleatoric uncertainty outside the scope of our analysis. Recent studies have shown that sources of uncertainty in deep learning can be disentangled using multi-step learning procedures, together with architectures designed to model aleatoric uncertainty; when coupled with methods such as SGLD, these approaches could enable uncertainty disentanglement.
In this respect, we find the preliminary work by Yi [yi2025cooperative] very promising. Although it is outside the scope of this paper, we believe that their approach could be used in conjunction with ours, and we plan to explore this synergy in future work.
@@ 6 Monte Carlo Study for Empirical Coverage
To estimate the performance of our model, we employ a Monte Carlo simulation study, similar to the one presented in [sarrias2018individual] for Latent Class discrete choice models. Here, we assume that the latent utility function follows the equation:
@@ 7 Revealed and Stated Preference Case Studies
In addition to the simulation study, we consider two case studies: one with revealed preference data and another with stated preference data. We use the revealed preference dataset constructed in [VILLARRAGA2025103132] for multinomial mode choice in NYC. For the stated preference data, we use the popular and publicly available [^4] Available in Apollo in R as the apollo_swissRouteChoiceData. dataset from a binary train choice experiment in Switzerland [vrtic2002impact].
@@ 7.1 Mode Choice in New York City
High income Indicator variable for high income level (>100>100k USD per year). $0.32$
@@ 8.1 Simulation Study
The empirical coverages for the MRS, illustrated in Figures 8 to 10 and summarized in Table 3, show that: i) our model achieves empirical coverage consistent with a 95% credibility level, regardless of dataset size; ii) the fully connected neural network fails to achieve adequate coverage for the smaller datasets but attains coverage close to 95% for the larger datasets; and iii) the model that ignores nonlinearities (the conditional logit model) achieves coverage close to 95% only for MRS associated with alternative attributes that have linear or near-linear relationships with the latent utilities, as we will show next [^6] Alternative attributes $x_{1}$ and $x_{2}$ have a linear relationship with the latent utilities, so the first MRS should be accurately estimated by a linear model..
@@ 8.2 Mode Choice in New York City
For the revealed preference case study using New York City multinomial mode choice data, we estimated values of travel time savings (VOTT) for train and car, as well as the out-of-sample accuracy of our model and a fully connected neural network. Standard discrete choice models, graph neural networks, and hierarchical Gaussian process models were previously estimated in [VILLARRAGA2025103132] and [2025arXiv250309786V].
@@
SOTA - 78.45  [^7] From the Skip-GNN presented by us in [2025arXiv250309786V].